seca+tana=p show that seca-tana=1/p ppppppppppppppllllllllllllllleeeeeeeeeeeeaaaaaaasssssssssseeeeeeeee help

seca+tana=p
show that seca-tana=1/p 
ppppppppppppppllllllllllllllleeeeeeeeeeeeaaaaaaasssssssssseeeeeeeee help

Grade:10

1 Answers

Vijay Mukati
8 years ago
Multiply the given relation by (seca – tana) both the sides and apply the formula (a-b)(a+b) = a2-b2..
On solving this, we will have sec2a – tan2a = p*(seca – tana) .
Now sincesec2a – tan2a = 1. Thereforeseca-tana=1/p.